a consumers income is $100. prices of goods X and Y are $1 per unit. suppose the consumer facing such prices chooses the bundle that includes 60 units of good Y. next, the price of good X increases to $2 per unit. the consumer's new choice involves a bundle with 40 units of good X. is X a normal good,an inferior good,or is there insufficient information to answer this question? explain using a graph and the concepts of income and substitution effects.?
Answer #1

Ques. Initially, Income Pe = = $100 $1 Budget Line Equation Pr. Qx & ly. ay = I 1.Q2 + . pQy = $100 x + ay = 100 Qxt to =100Since, we can see that as price of good X increases from $1 per unit to $2 per unit the demand will remain at 40 units. Since only price information is given nothing is mentioned that How income will change as price changes and also Utilitu Function is not given we can not say whether the good is inferior or normal.

Hence there is insufficient information to answer this.

& Pecice Effect is the son of Substitution Effect and Income effect Income effect o Adjusting Puschasing Power, by keeping reI The magnitude of both will docide the Price Effect PE = SE + LE = Ive + I tue love) Initially, ŚË > IE as o resueet - Pace@ In case guitial Budget dire Normeer 62 - IC final Budget Live Compensa keel Budget x x 122 B Price Effect XX, = SE X,X2 = IAs you can here we need Indifference curve to show substitution effect and income effect.